(a) Write the claim mathematically and identify Upper H 0 and Upper H Subscript a. ?(b) Find the critical? value(s) and identify the rejection? region(s). (c) Find the standardized test statistic.? (d) Decide whether to reject or fail to reject the null hypothesis.

In a sample of 1000 home? buyers, you find that 473 home buyers found their real estate agent through a friend. At a=0.02?, can you reject the claim that 50% of home buyers find their real estate agent through a? friend?
Answer #1

Sample proportion \hat{p} = 473 / 1000 = 0.473

a)

H0: p = 0.50

Ha: p \neq 0.50

b)

Critical values at 0.02 level are -2.326, 2.326

Rejecton region is reject H0 if test statistics < -2.326 OR test statistics > 2.326

c)

Test statistics

Z = \hat{p} - p / sqrt( p ( 1 - p) / n)

= 0.473 - 0.50 / sqrt( 0.50 * 0.50 / 1000)

= -1.708

This is test statistics value.

d)

Since test statistics value falls in non-rejection region, we do not have sufficient evidence

to reject H0.

We conclude at 0.02 level that we fail to support the claim.